OpenStudy (anonymous):

What is the equation, in point-slope form, of the line that has a slope of 6 and passes through the point (–1, –8)?

OpenStudy (anonymous):

Eq of line in slope point form is given as: \[y-y_1= m(x-x_1) \rightarrow y-(-8) = 6[x-(-1)]\] y+8=6(x+1) y+8=6x+6 y= 6x-2 is the required eq in slope point form.

undeadknight26 (undeadknight26):

*gasp* im undead too...

OpenStudy (anonymous):

yah well my i have a reason for the undead name

undeadknight26 (undeadknight26):

i do too!

OpenStudy (anonymous):

if you dont mind me asking what is it lol

undeadknight26 (undeadknight26):

I neva said it was a good reason XD im a gamer and i liked the name...so i took up the name as undeadknight26!

OpenStudy (anonymous):

yah well im part of something called undead army lol